OpenStudy (jenniferjuice):

Solve: 4y − y + 11 = 38

OpenStudy (bakonloverk):

combine your like terms and then solve the equation

OpenStudy (anonymous):

yeah so subtract 4y from 1y

OpenStudy (jenniferjuice):

4y-y+11=38 4y-y=38-11 4y-y=27 4y-1y=27 3y=27 3y/3=27/3 y=9

OpenStudy (bakonloverk):

then do 38 - 11 and with that answer divide it by the answer of 4y-y

OpenStudy (jenniferjuice):

?

OpenStudy (bakonloverk):

yes thats right

OpenStudy (anonymous):

yup you go it! good job

OpenStudy (jenniferjuice):

thamkyou :x

OpenStudy (bakonloverk):

np